# Flat band effects on the ground-state BCS-BEC crossover in atomic Fermi gases in a quasi-two-dimensional Lieb lattice

### v2

@@ Abstract
Figure 1: $\Delta$ as a function of $-g$ for various $n$ in the weak coupling BCS regime, with $\mu$ close to or inside the flat band. Shown in the inset is the DOS for strict 2D case, where the (red) arrow denotes the flat band.
@@
We first investigate the flat band effects on the pairing gap behavior, when the Fermi level is close to or within the flat band. Due to the particle-hole symmetry, we restrict $n\leq 3$, and the Fermi level enters the flat band for $n\geq 2$ in the noninteracting limit. Plotted in Fig. 1 is $\Delta$ versus $-g$ (in units of $t$) for various densities near $n\lesssim 2$ and $n\in[2,3]$ in the weak coupling BCS regime. Shown in the inset is the DOS for the strict 2D case, where the (red) vertical arrow denotes the flat band, along with two VHS’s in the upper and lower bands, corresponding to $n=1$ and $5$, respectively. For $n<2$, including $n=1.95$, $\Delta$ exhibits an exponentially activated behavior in the weak coupling regime, similar to its counterpart in 3D continuum and 3D cubic lattices. This is expected in BCS theory, under the approximation of a constant DOS near the Fermi surface. However, as the density goes higher than $n=2$, the Fermi level enters the flat band with the lower band fully filled, and the behavior of $\Delta$ changes into an unusual power law, which can be attributed to the breakdown of the constant DOS approximation. Similar behavior has also been predicted within the dynamical mean-field theory in a quasi 2D repulsive Lieb lattice, where the magnetism as a function of the interaction changes from exponential to power law behavior at half filling [31]. The behavior of $\Delta$ over a broad range is shown in the inset of Fig. 2(b) on a log-log scale for a lower density, $n=0.6$, which shows that the numerical result (black solid line) approaches its analytical BEC asymptote (red dashed line) in the deep BEC regime, as given by Eq. (9).
@@
This unusual power-law behavior of $\Delta$ versus $g$ can be explained following Ref. [31], despite the different signs of the interaction. Using the normalized, dimensionless notation $\overline{X}\equiv X/W$, where $W=4\sqrt{2}t+4t_{z}$ is the bandwidth, the rescaled, dimensionless DOS can be approximated as $\overline{\rho}(\overline{\varepsilon})\equiv W\rho(\overline{\varepsilon})=4\theta(1/2-|\overline{\varepsilon}-1/2|)+2\delta(\overline{\varepsilon}-1/2)$, where the $\delta$ function represents the flat band.
For $n>2$, using this $\overline{\rho}(\overline{\varepsilon})$ for the integral in the gap equation (2), we obtain
$\overline{\Delta}=\frac{n-2}{4\mathcal{W}(\frac{n-2}{4}\exp(\frac{1}{2|\overline{g}|}))}\,,$
@@
where $\mathcal{W}$ is the Lambert $\mathcal{W}$-function, which is the inverse function of $f(\mathcal{W})=\mathcal{W}\exp(\mathcal{W})$. For small $x\rightarrow 0$, $\mathcal{W}(x)\approx x-x^{2}+3x^{3}/2$. This leads to $\overline{\Delta}\propto(n-2)|\overline{g}|$ for $n>2$ in the weak coupling regime. At low $n<2$, the $\delta$-function term in $\overline{\rho}(\overline{\varepsilon})$ becomes irrelevant, so that we recover the ordinary BCS behavior, $\overline{\Delta}\propto\exp(1/n\overline{g})$.
@@
Figure 2: Evolution of $\mu$ as a function of $-g/t$ for small $0.6\leq n\leq 1.2$ in panel (a) and large $1.9\leq n\leq 2.7$ in panel (b), as labeled. Shown in the insets of panels (a) and (b) is a comparison between the full numerical solution (black solid) and the analytical BEC asymptotic behavior (red dashed) of $-\mu$ and $\Delta$, respectively, as a function of $-g/t$ on a log-log scale for $n=0.6$.
@@
Next, we show in Fig. 2 the behaviors of $\mu$ as a function of $-g$ for various $n$, so that the Fermi level changes from (a) the lower band below ($n=0.6$) and around the VHS ($0.9\leq n\leq 1.2$) to (b) near ($n=1.9,1.95$) or inside the flat band ($n=2,2.3,2.7$). As shown in the inset of Fig. 2(a), $-\mu$ (black solid line, for $n=0.6$) approaches its BEC asymptotic behavior (red dashed line) for $-g/t>10$. For $n\leq 1$ in Fig. 2(a), $\mu$ decreases monotonically as the pairing interaction becomes stronger, similar to that in a regular one-band model below half filling. However, for $n=1.1,1.2$, $\mu$ exhibits a remarkable nonmonotonic behavior in the weak coupling BCS regime; $\mu$ increases first as the pairing becomes stronger, and then starts to decrease after passing a maximum. Such nonmonotonicity is also found in a 2D optical lattice with a strong lattice effect, which is comprised of two lattice and one continuum dimensions [41, 42]. In a quasi-2D Lieb lattice, the lower band has two VHS’s, at $\varepsilon=(2\sqrt{2}-2)t\approx 0.8284t$ and $\varepsilon=(2\sqrt{2}-2)t+4t_{z}\approx 0.8684t$. For $n=1.1,1.2$, $\mu>0.8684t$ for small $|g|$, i.e., the Fermi level sits slightly above the VHS’s, where the DOS $\rho(\varepsilon)$ has a negative slope, so that the pairing becomes hole-like, for which $\mu$ increases as the pairing becomes stronger, as in a 3D cubic lattice above half filling. Indeed, in the weak coupling limit, there is generally an approximate particle-hole symmetry at the VHS, as the DOS in the vicinity of the VHS dominates. As pairing strength increases further, the contribution of the DOS below the VHS’s comes in, so that $\mu$ starts to decrease again. As $n$ increases and approaches $n=2$, the flat band gradually affects the behaviors of $\mu$ in the BCS regime. For $n=1.9,1.95$ in Fig. 2(b), $\mu$ remains nearly a constant before decreasing as $-g$ increases in the BCS regime. For $2\leq n\leq 3$, $\mu$ enters the flat band, and starts to decrease from roughly the same noninteracting limit $\mu_{0}\approx 2\sqrt{2}t$. While the nearly constant $\mu$ for $n=1.9,1.95$ is in accord with the exponentially small gap in the BCS regime (Fig. 1), a power-law decrease in $\mu$ can be readily seen for $n\in[2,3]$, commensurate with the power-law increase of $\Delta$ as a function of $-g$.
@@
Shown in Fig. 3 are the in-plane (left column) and out-of-plane (right column) superfluid density, from top to bottom, for $n=0.6$, $2$ and $2.7$, respectively, as well as the conventional $(n_{s}/m)^{\rm conv}_{\parallel}$ (blue) and geometric part $(n_{s}/m)^{\rm geom}_{\parallel}$ (red curves) of $(n_{s}/m)_{\parallel}$. For $n=0.6$, both the in-plane $(n_{s}/m)_{\parallel}$ and out-of-plane $(n_{s}/m)_{z}$, as well as the conventional $(n_{s}/m)^{\rm conv}_{\parallel}$, exhibit a roughly constant behavior as a function of $-g$ in the weak coupling regime, similar to the constant superfluid density $n/m$ in 3D free space. On the contrary, the geometric part $(n_{s}/m)^{\rm geom}_{\parallel}$ vanishes in the noninteracting limit. Commensurate with the evolution of $\Delta$ versus $-g$, as $n$ increases to $n\geq 2$, the behaviors of both $(n_{s}/m)^{\rm conv}_{\parallel}$ and $(n_{s}/m)^{\rm geom}_{\parallel}$, as well as $(n_{s}/m)_{\parallel}$, evolve into power laws as a function of $-g$. More importantly, $(n_{s}/m)^{\rm conv}_{\parallel}$ now approaches zero in the noninteracting limit, unlike in a simple cubic lattice. This can be attributed to a few reasons. (i) Without the out-of-plane contribution, the DOS from the upper and lower band is zero when $\mu$ falls inside the flat band. (ii) There is a perfect particle-hole symmetry at this $\mu$ so that the contributions from particles and holes cancel out in $(n_{s}/m)_{\parallel}$; the pairing is neither particle-like nor hole-like. (iii) The quantum geometric contribution relies on a finite gap $\Delta$, and hence vanishes in the zero $g$ limit. (iv) The flat band does not contribute to superfluidity without the quantum geometric effect. As $\mu$ moves away from the flat band with increasing $|g|$, the particle-hole cancellation breaks down, leading to a rising $(n_{s}/m)_{\parallel}$. As $|g|$ increases further toward the BEC regime, $(n_{s}/m)_{\parallel}$ starts to decrease due to the lattice effects. Without a geometric contribution, the out-of-plane superfluid density $(n_{s}/m)_{z}$ exhibits a behavior similar to that of $(n_{s}/m)^{\rm conv}_{\parallel}$ in the weak coupling regime for $n=0.6$ and $2$. However, difference appears as $\mu$ enters the flat band. As shown in the inset of Fig. 3(f), when $n$ becomes slightly higher than 2, $(n_{s}/m)_{z}$ starts to increase as the noninteracting limit is approached. At $n=2.7$, shown in Fig. 3(f), this increase is so dramatic that $(n_{s}/m)_{z}$ becomes monotonically decreasing with $|g|$. This distinct behavior of $(n_{s}/m)_{z}$, compared to $(n_{s}/m)^{\rm conv}_{\parallel}$, results from the broadening of the flat band due to the small out-of-plane dispersion with $t_{z}/t=0.01$.
@@
Figure 4: $\kappa$ versus $-g$ for (a) small $n$ below ($n=0.6$) and close to the van Hove singularities around $n=1$ and (b) large $n$ near the bottom of the flat bat $n=2$. Shown in the inset of panel (a) is the comparison between the full numerical solution (black solid) and the analytical BEC asymptotic behavior (red dashed) of $\kappa$ as a function of $-g$ on a log-log scale for $n=0.6$.
@@
We now plot in Fig. 4 the compressibility $\kappa$ versus $-g$ for (a) small $n$ below and around the VHS and (b) large $n$ near the flat band bottom. Figure 4(a) indicates that the noninteracting value of $\kappa$ reaches a local maximum at the VHS $n=1$ as a function of $n$, since this value is given by the DOS. Furthermore, for $n\in[0.9,1.2]$, $\kappa$ exhibits a nonmonotonic dependence on $|g|$, which is largely related to the nonmonotonic behavior of $\mu(g)$. As the interaction goes stronger into the BEC regime, all fermions pair up and the two-body binding energy of the fermions dominates $\mu$. As a consequence, $\kappa$ decreases and approaches the same $n$-independent BEC asymptote. Indeed, as shown in the inset for $n=0.6$, $\kappa$ approaches nicely its analytical $n$-independent BEC asymptote when $-g/t>10$. The nearly constant $\kappa$ for $n=1.9$ and 1.95 in Fig. 4(b) is clearly associated with the constant behavior of $\mu$ in the BCS regime shown in Fig. 2(b). For $n\geq 2$, $\kappa$ increases sharply as the interaction strength decreases toward zero. Here the Fermi level falls within the flat band, and thus senses a huge DOS, which is equal to the noninteracting $\kappa$. Therefore, there is a jump of $\kappa(g=0)$ as $n$ crosses the $n=2$ boundary, as shown in Fig. 4(b).
@@
Figure 5: The ground-state phase diagram in the $n-g$ plane. The black dot-dashed $\mu=0$ curve divides the plane into fermionic and bosonic superfluid regimes. Enclosed within the (red) $B=0$ and (blue) $B_{z}=0$ lines is a PDW ground state (gray shaded region).
@@
Finally, we present in Fig. 5 the ground-state phase diagram in the $n$$g$ plane. Here the positivity of $\xi^{2}=a_{0}B$ and $\xi_{z}^{2}=a_{0}B_{z}^{2}$ constitutes two stability conditions for the superfluid phase. The (black dot-dashed) $\mu=0$ curve separates the fermionic superfluid regime on the upper left from the bosonic superfluid regime on the lower right. A PDW ground state with negative $\xi^{2}<0$ and/or $\xi_{z}^{2}<0$ emerges in the grey shaded region, enclosed inside the red $B=0$ curve and blue $B_{z}=0$ curve. Note that the upper branch of $B=0$ overlaps nearly precisely with $B_{z}=0$, and both are close to but falls slightly inside the fermionic side of the $\mu=0$ line. The PDW state at intermediate and strong coupling for relatively large $n$ is associated with the strong repulsive inter-pair interaction $U$ and relatively low kinetic energy of the pairs, $T$, which leads naturally to Wigner crystallization. Indeed, for $N$ pairs, the inter-pair interaction energy $E_{P}$ scales as $N^{2}U/2$ whereas the kinetic energy $E_{K}$ scales as $NT$, allowing $E_{P}>E_{K}$ with a large $U$ and lattice suppressed $T$. Technically, the sign of $\xi^{2}$ (or $\xi_{z}^{2}$) becomes negative between the $B=0$ (or $B_{z}=0$) curve, indicating that the minimum of the pair dispersion $\tilde{\Omega}_{\textbf{q}}$ has moved from $\mathbf{q}=0$ to a finite $\mathbf{q}$, with the crystallization wave vector in the $xy$ plane (or in the $z$ direction). The PDW state has been observed in experiments [50], but it is still unclear whether it can sustain superfluidity and thus becomes a supersolid, which will be left to a future study. Similar PDW ground state has been found in 3D optical lattice at high density [39], 2D optical lattice with strong lattice effect [41, 42], optical lattice in mixed dimensions [51], and in dipolar Fermi gases [40].

### v3

@@ III Numerical Results and Discussions
Figure 2: $\Delta$ as a function of $-g$ for various $n$ in the weak coupling BCS regime, with $\mu$ close to or inside the flat band. Shown in the inset is the DOS for strict 2D, where the (red) arrow denotes the flat band.
@@
We first investigate the flat band effects on the pairing gap behavior, when the Fermi level is close to or within the flat band. Due to the particle-hole symmetry, we restrict ourselves to $n\leq 3$, and the Fermi level enters the flat band for $n\geq 2$ in the noninteracting limit. Plotted in Fig. 2 is $\Delta$ versus $-g$ (in units of $t$) for various densities near $n\lesssim 2$ and $n\in[2,3]$ in the weak coupling BCS regime. Shown in the inset is the DOS for the strict 2D case, where the (red) vertical arrow denotes the flat band, along with two VHS’s in the upper and lower bands, corresponding to $n=1$ and $5$, respectively. For $n<2$, including $n=1.95$, $\Delta$ exhibits an exponentially activated behavior in the weak coupling regime, similar to that in 3D continuum and 3D cubic lattices. This is expected in BCS theory, assuming a roughly constant DOS near the Fermi surface. However, as the density $n$ goes above $2$, the Fermi level enters the flat band with the lower band fully filled, and the gap $\Delta(g)$ as a function of the interaction exhibits an unusual power law, which can be attributed to the breakdown of the constant DOS approximation. Similar behavior has also been predicted within the dynamical mean-field theory in a quasi 2D repulsive Lieb lattice, where the magnetism as a function of the interaction $g$ changes from exponential to power law behavior at half filling [31]. The behavior of $\Delta$ over a broad range is shown in the inset of Fig. 3(b) on a log-log scale for a lower density, $n=0.6$. Indeed, the numerical result (black solid line) approaches its analytical BEC asymptote (red dashed line) in the deep BEC regime, as given by Eq. (8).
@@
This unusual power-law behavior of $\Delta$ versus $g$ can be explained following Ref. [31], despite the different signs of the interaction. Using the normalized, dimensionless notation $\overline{X}\equiv X/W$, where $W=4\sqrt{2}t+4t_{z}$ is the bandwidth, the rescaled, dimensionless DOS can be simplified as $\overline{\rho}(\overline{\varepsilon})\equiv W\rho(\overline{\varepsilon})=4\theta(1/2-|\overline{\varepsilon}-1/2|)+2\delta(\overline{\varepsilon}-1/2)$, where the $\delta$ function represents the flat band.
For $n>2$, using $\overline{\rho}(\overline{\varepsilon})$ in the integral in the gap equation (2), we obtain from the gap equation (2)
$\overline{\Delta}=\frac{n-2}{8}/\mathcal{W}(\frac{n-2}{8}\exp(\frac{1}{4|\overline{g}|})),$
@@
where $\mathcal{W}$ is the Lambert $\mathcal{W}$-function, which is the inverse function of $f(\mathcal{W})=\mathcal{W}\exp(\mathcal{W})$ (See Appendix B for a derivation). For $x\gg 1$, $\mathcal{W}(x)\approx\ln(x)-\ln(\ln(x))$. To leading order, this yields $\overline{\Delta}\approx\frac{n-2}{2}|\overline{g}|$ for weak interactions. At low $n<2$, the $\delta$-function term in $\overline{\rho}(\overline{\varepsilon})$ becomes irrelevant, so that we recover the ordinary exponential BCS behavior.
@@
Figure 3: Evolution of $\mu$ as a function of $-g/t$ for (a) small $0.6\leq n\leq 1.2$ and (b) large $1.9\leq n\leq 2.7$, as labeled. In the insets, the full numerical solution (black solid) for $n=0.6$ is compared with the analytical BEC asymptotic behavior (red dashed curves) for $-\mu$ and $\Delta$, respectively, as a function of $-g/t$ on a log-log scale.
@@
Next we show in Fig. 3 the behaviors of $\mu$ as a function of $g$ for various densities $n$, so that the Fermi level changes from (a) the lower band below ($n=0.6$) and around the VHS ($0.9\leq n\leq 1.2$) to (b) near ($n=1.9,1.95$) or inside the flat band ($n=2,2.3,2.7$). As shown in the inset of Fig. 3(a), $-\mu$ (black solid line, for $n=0.6$) approaches its BEC asymptotic behavior (red dashed line) for $-g/t>10$. For $n\leq 1$ in Fig. 3(a), the chemical potential $\mu$ decreases monotonically as the pairing interaction becomes stronger, similar to that in a regular one-band model below half filling. However, for $n=1.1,1.2$, $\mu$ becomes remarkably nonmonotonic in the weak coupling regime; $\mu$ increases first with $|g|$, and then starts to decrease after passing a maximum. Such nonmonotonicity is also found in a 2D optical lattice with a strong lattice effect, which is comprised of two lattice and one continuum dimensions [41, 42]. In a quasi-2D Lieb lattice, the lower band has two VHS’s, at $\varepsilon=(2\sqrt{2}-2)t\approx 0.8284t$ and $\varepsilon=(2\sqrt{2}-2)t+4t_{z}\approx 0.8684t$. For $n=1.1,1.2$, $\mu>0.8684t$ for small $|g|$, i.e., the Fermi level sits slightly above the VHS’s for small $|g|$, where the DOS $\rho(\varepsilon)$ has a negative slope, so that the pairing becomes hole-like, for which $\mu$ increases as the pairing becomes stronger, as in a 3D cubic lattice above half filling. Indeed, in the weak coupling limit, there is generally an approximate particle-hole symmetry at the VHS, as the DOS in the vicinity of the VHS’s dominates. As $|g|$ increases further, the contribution of the DOS below the VHS’s comes in, so that $\mu$ starts to decrease again. As $n$ increases and approaches $n=2$, the flat band gradually affects the behaviors of $\mu$ in the BCS regime. For $n=1.9,1.95$ in Fig. 3(b), $\mu$ remains nearly a constant before decreasing with $|g|$ in the BCS regime. For $2\leq n\leq 3$, $\mu$ enters the flat band, and starts to decrease from roughly the same noninteracting limit $\mu_{0}\approx 2\sqrt{2}t$. While the nearly constant $\mu$ for $n=1.9,1.95$ is in accord with the exponentially small gap in the BCS regime (Fig. 2), a power-law decrease in $\mu$ can be readily seen for $n\in[2,3]$, commensurate with the power-law increase of $\Delta$ as a function of $-g$.
@@
Shown in Fig. 4 are the in-plane (left column) and out-of-plane (right column) superfluid density, from top to bottom, for $n=0.6$, $2$ and $2.7$, respectively, as well as the conventional $(n_{s}/m)^{\rm conv}_{\parallel}$ (blue) and geometric part $(n_{s}/m)^{\rm geom}_{\parallel}$ (red curves) of $(n_{s}/m)_{\parallel}$. For $n=0.6$, both the in-plane $(n_{s}/m)_{\parallel}$ and the out-of-plane $(n_{s}/m)_{z}$, as well as the conventional $(n_{s}/m)^{\rm conv}_{\parallel}$, are roughly constant for weak coupling, similar to the constant superfluid density $n_{s}/m=n/m$ in 3D continuum. On the contrary, the geometric part $(n_{s}/m)^{\rm geom}_{\parallel}$ vanishes in the noninteracting limit $g\rightarrow 0$. Commensurate with the evolution of $\Delta$ versus $-g$, as $n$ increases to $n\geq 2$, the behaviors of both $(n_{s}/m)^{\rm conv}_{\parallel}$ and $(n_{s}/m)^{\rm geom}_{\parallel}$, as well as $(n_{s}/m)_{\parallel}$, evolve into power laws as a function of $-g$. More importantly, $(n_{s}/m)^{\rm conv}_{\parallel}$ now approaches zero as $g\rightarrow 0$, unlike in a simple cubic lattice. This can be attributed to a few reasons. (i) Without the out-of-plane contribution, the DOS from the upper and lower band is zero when $\mu$ falls inside the flat band. (ii) There is a perfect particle-hole symmetry at this $\mu$ so that the contributions from particles and holes cancel out in $(n_{s}/m)_{\parallel}$; the pairing is neither particle-like nor hole-like. (iii) The quantum geometric contribution relies on a finite gap $\Delta$, and hence vanishes in the zero $g$ limit. (iv) The flat band does not contribute to superfluidity without the quantum geometric effect. As $\mu$ moves away from the flat band with increasing $|g|$, the particle-hole cancellation breaks down, leading to a rising $(n_{s}/m)_{\parallel}$. As $|g|$ increases further toward the BEC regime, $(n_{s}/m)_{\parallel}$ starts to decrease due to the lattice effects. Without a geometric contribution, the out-of-plane superfluid density $(n_{s}/m)_{z}$ exhibits a behavior similar to that of $(n_{s}/m)^{\rm conv}_{\parallel}$ in the weak coupling regime for $n=0.6$ and $2$. However, difference appears as $\mu$ enters the flat band. As shown in the inset of Fig. 4(f), when $n$ becomes slightly higher than 2, $(n_{s}/m)_{z}$ starts to increase as $g\rightarrow 0$. At $n=2.7$, shown in Fig. 4(f), this increase is so dramatic that $(n_{s}/m)_{z}$ becomes monotonically decreasing with $|g|$. This distinct behavior of $(n_{s}/m)_{z}$, compared to $(n_{s}/m)^{\rm conv}_{\parallel}$, results from the broadening of the flat band due to the small out-of-plane dispersion.
@@
Figure 5: $\kappa$ versus $-g$ for (a) small $n$ below ($n=0.6$) and close to the VHS’s around $n=1$ and (b) large $n\simeq 2$ near the flat band bottom. Shown in the inset is the full numerical solution of $\kappa(g)$ (black solid), compared with its BEC asymptotic expression $\kappa=-2/g$ (red dashed curve) for $n=0.6$ on a log-log scale.
@@
Plotted in Fig. 5 is the compressibility $\kappa$ versus $g$ for (a) small $n$ below and around the VHS and (b) large $n$ near and in the flat band. Figure 5(a) indicates that the noninteracting value of $\kappa$ reaches a local maximum at the VHS $n=1$ as a function of $n$, since this value is given by the DOS. Furthermore, for $n\in[0.9,1.2]$, $\kappa$ varies nonmonotonically with $g$, largely related to the nonmonotonic behavior of $\mu(g)$. As $|g|$ increases into the BEC regime, all fermions pair up and the two-body binding energy dominates $\mu$. Consequently, $\kappa$ decreases and approaches the same $n$-independent BEC asymptote. Indeed, as shown in the inset for $n=0.6$, $\kappa$ approaches nicely the $n$-independent analytical BEC asymptote when $-g/t>10$. The nearly constant $\kappa$ for $n=1.9$ and 1.95 in Fig. 5(b) is clearly associated with the constant behavior of $\mu$ in the BCS regime shown in Fig. 3(b). For $n\geq 2$, $\kappa$ increases sharply as $g\rightarrow 0$. Here the Fermi level falls within the flat band, and thus the noninteracting $\kappa$ is given by the huge DOS. Therefore, there is a jump of $\kappa(g=0)$ as $n$ crosses the $n=2$ boundary, as shown in Fig. 5(b).
@@
Figure 6: Ground-state phase diagram in the $n-g$ plane. The (black dot-dashed) $\mu=0$ curve divides the plane into fermionic and bosonic superfluid regimes. Enclosed within the (red) $B=0$ and (blue) $B_{z}=0$ lines is a PDW ground state (gray shaded region).
@@
Finally, we present in Fig. 6 the ground-state phase diagram in the $n$$g$ plane. Here the positivity of $\xi^{2}=a_{0}B$ and $\xi_{z}^{2}=a_{0}B_{z}^{2}$ constitutes two stability conditions for the superfluid phase. The (black dot-dashed) $\mu=0$ curve separates the fermionic superfluid regime on the upper left from the bosonic superfluid regime on the lower right. A PDW ground state with negative $\xi^{2}<0$ and/or $\xi_{z}^{2}<0$ emerges in the grey shaded region, enclosed inside the (red) $B=0$ and (blue) $B_{z}=0$ curves. Note that the upper branch of $B=0$ overlaps nearly precisely with $B_{z}=0$, and both are close to but falls slightly inside the fermionic side of the $\mu=0$ line. The PDW state at intermediate and strong coupling for relatively large $n$ is associated with the strong repulsive inter-pair interaction $U$ and relatively low kinetic energy of the pairs, $T$, which leads naturally to Wigner crystallization. Indeed, for $N$ pairs, the inter-pair interaction energy $E_{P}$ scales as $N^{2}U/2$ whereas the kinetic energy $E_{K}$ scales as $NT$, allowing $E_{P}>E_{K}$ with a large $U$ and lattice suppressed $T$. Technically, the sign of $\xi^{2}$ (or $\xi_{z}^{2}$) becomes negative between the $B=0$ (or $B_{z}=0$) branches, where the pair dispersion $\tilde{\Omega}_{\textbf{q}}$ reaches a minimum at a finite $\mathbf{q}\neq 0$, with the crystallization wave vector in the $xy$ plane (or in the $z$ direction). The PDW state has been observed in experiments [52], but it is still unclear whether it can sustain superfluidity and thus becomes a supersolid, which will be left to a future study. A similar PDW ground state has been found in 3D lattices at high density [39], 2D optical lattices with strong lattice effect [41, 42], optical lattices in mixed dimensions [53], and in dipolar Fermi gases [40].
IV Conclusions